OpenStudy (anonymous):

If sin A= radical five over three, then cos 2A equals

OpenStudy (anonymous):

Using the trig identity: cos(2A) = cos²A − sin²A = 1 − 2 sin²A , we can plug in what we already know: 1 - 2 ((sqrt 5)/3)^2 = 1- 2(5/9) = 1- 10/9 = -1/9